OpenStudy (anonymous):

Can Someone please help! Find a particular solution yp of the differential equation 6y''+6y'+8y=(cos^2(x)) using the Method of Undetermined Coefficients. Primes denote derivatives with respect to x.

OpenStudy (anonymous):

We can use this: cos^2(x)=[1+cos(2x)]/2 So we are looking for a yp like this: yp=A+B*cos(2x)+C*sin(2x) Now we find the derivatives: yp'=-2B*sin(2x)+2C*cos(2x) yp''=-4B*cos(2x)-4C*sin(2x) Now we plug that into the equation: 6*(-4B*cos(2x)-4C*sin(2x))+6(-2B*sin(2x)+2C*cos(2x))+8*(A+B*cos(2x)+C*sin(2x))= =cos(2x)*[-24B+12C+8B] +sin(2x)*[-24C-12B+8C]+8A =[1+cos(2x)]/2 So then we get this: 8A=1/2 -24C-12B+8C=0 -24B+12C+8B=1/2 A=1/16 Now we have to solve for B and C

OpenStudy (anonymous):

-12B+16C=0 -16B+12C=1/2 I tihnk that you can solve this, right?

OpenStudy (anonymous):

so you'd get B=0 and C=(1/24) right

OpenStudy (anonymous):

but then that would be incorrect for the end solution

OpenStudy (anonymous):

I got B=-1/14 C=-3/56

OpenStudy (anonymous):

So then the solution is: yp=1/8-(1/14)*cos(2x)-(3/56)*sin(2x)

OpenStudy (anonymous):

How did you solve for B?

OpenStudy (anonymous):

We can solve like this: -12B+16C=0 -16B+12C=1/2 Solve for B on the first one. B=(16/12)*C then plug that into the second one: -16*(16C/12)+12C=1/2 There we can solve for C: -256C/12+12C=1/2 (144-256)C/12=1/2 -112C/12=1/2 C=-12/(2*112) C=-3/56 Now we go back to the first one: -12B+16*(-3/56)=0 And now we solve for B -12B=48/56 B=-1/14

OpenStudy (anonymous):

ok thank you so much!!

OpenStudy (anonymous):

Thanks to you! You are welcome! :)

OpenStudy (anonymous):

wait shouldnt the solution be 1/16-(1/14)*cos(2x)-(3/56)*sin(2x)

OpenStudy (anonymous):

Yes, you are right! :) Sorry about my confusion.

OpenStudy (anonymous):

unfortunately thats still wrong! It's part of thisonline web homework and its saying its wrong

OpenStudy (anonymous):

A sec please.

OpenStudy (anonymous):

-24C-12B+8C=0 -24B+12C+8B=1/2 I found my mystake: The equation that we have to solve is this one: -16C-12B=0 -16B+12C=1/2 B=-1/50 C=3/200 Sorry.

OpenStudy (anonymous):

and A is 1/16?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Is that correct now?

OpenStudy (anonymous):

yes it is! thanks a bunch!!!

OpenStudy (anonymous):

You are welcome! :)

OpenStudy (anonymous):

:D